Molly has received scores of 85, 88, 87, and 90 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 87?
Answer by Alan3354(69443)

Molly has received scores of 85, 88, 87, and 90 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 87?
------------
T = 5th score
---
(85+88+87+90+ T)/5 = 87*5 for average of 87
350 + T = 435
T = 85
--> T >= 85